# AWM[Uncaptioned image]: Accurate Weight-Matrix Fingerprint for Large Language Models

### v1

@@ 6 Experiments
Baselines. We compare our method against two advanced baselines: the weight-based method HuRef (Zeng et al., 2024), and the representation-based method REEF (Zhang et al., 2025).
@@ 6.3 Robustness Verification on 60 Offspring Models
Table 2: Absolute Z-scores ($\uparrow$) of positive samples for HuRef, REEF, and our method. The full names of abbreviations are provided in Appendix C.
@@ 6.4 Overall Evaluation
To synthesize our findings, we conducted a comprehensive evaluation against HuRef and REEF using the full dataset of 60 positive and 90 negative model pairs.
@@ 7 Conclusions
Acknowledgement
This work is sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402) and the National Natural Science Foundation of China (NSFC) grant (No.62576211).
Appendix A Definitions and Notations
A.1 LLM Architecture
@@
$\text{RMSNorm}(X,\omega_{A}^{(l)},\epsilon_{A})=\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon_{A}\mathbf{1}_{m})\cdot X\cdot\text{diag}(\omega_{A}^{(l)}).$ (6)
Definition A.2 (Self-Attention with RoPE).
@@
Here $x_{i}^{\top}\in\mathbb{R}^{n}$ is the i-th row of $X$, $\theta$ is the frequency base for RoPE and $R_{\theta}$ is the rotation matrix (Su et al., 2024).
RoPE’s definition in Section A.1 is slightly different from their implementations, but still an equivalent version to Su et al. (2024).
Definition A.3 (Feed-forward Networks).
@@
We abuse the notations in Section A.1 by denoting $\text{RMSNorm}_{\text{attn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})$ as the attention norm and $\text{RMSNorm}_{\text{ffn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{ffn}}^{(l)},\epsilon_{A})$ the ffn norm at layer $l$ of model A. Then, the definition of a Transformer block at layer $l$ of model A can be summartized as follows.
Definition A.4 (Transformer Block).
@@
A.2 Central Kernel Alignment (CKA)
@@
where $\text{HSIC}(K_{X_{1}},K_{X_{2}})=\frac{1}{(m-1)^{2}}\text{tr}(K_{X_{1}}H_{m}K_{X_{2}}H_{m})$, $(K_{X_{1}})_{ij}=k((X_{1})_{i},(X_{1})_{j})$, $(K_{X_{2}})_{ij}=k((X_{2})_{i},(X_{2})_{j})$ are kernel matrices with the kernel function $k(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}_{+}$, and $H_{m}=I_{m}-\frac{1}{m}\mathrm{1_{m}}\mathrm{1_{m}}^{\top}$ is the centering matrix. The linear kernels give $K_{X}=XX^{\top}$ for a matrix $X$. To reduce the finite-sample bias of Eq. (6), we use the unbiased HSIC estimator (Song et al., 2007). Let $\tilde{K}_{X_{i}}$ be $K_{X_{i}}$ with its diagonal set to zero, i.e., $(\tilde{K}_{X_{i}})_{ii}=0$. With $\mathrm{1}_{m}$ the all-ones vector of length $m$, the unbiased HSIC is
@@
Appendix B Proofs and Discussions
B.1 Proof for Theorem 3.1
We give the proof based on the definition of CKA with linear kernels in Appendix A.2. Given input matrices $X_{1},X_{2}$, orthogonal matrices $U_{1},U_{2}$ and constants $c_{1},c_{2}$, the kernel functions yield
@@ Appendix B Proofs and Discussions
B.3 Proof for Theorem 4.3 and Discussions on Other Input Manipulations
We first show a proof for the claims in Theorem 4.3. The core lies in preserving the diagonal structure of RMSNorm parameters in Section A.1. Given the input $X$ to layer $l$, the manipulations change it to $cXPD$ and fed the manipulated input into the attention norm of model B. In order to recover the manipulations on inputs after RMSNorm, i.e. for any manipulation $T$ in the input $X$ (an example is $T(X)=cXPD$),
@@
$=\text{diag}^{-\frac{1}{2}}\left((cXPD)(cXPD)^{\top}\mathbf{1}_{m}+\epsilon_{B}\mathbf{1}_{m}\right)\cdot cXPD\cdot\text{diag}(\omega_{B,\text{attn}}^{(l)})$
$=\text{diag}^{-\frac{1}{2}}(c^{2}XX^{\top}\mathbf{1}_{m}+c^{2}\epsilon_{A}\mathbf{1}_{m})\cdot cXPD\cdot cD^{\top}P^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})PD$
$=c\cdot\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon_{A}\mathbf{1}_{m})\cdot X\cdot\text{diag}(\omega_{A,\text{attn}}^{(l)})\cdot PD$
@@
However,
$\text{RMSNorm}(XU,\omega,\epsilon)=\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon\mathbf{1}_{m})\cdot XU\cdot\text{diag}(\omega^{(l)}).$
@@
$\text{RMSNorm}(XM,\omega^{(l)},\epsilon)=\text{diag}^{-\frac{1}{2}}(XMM^{\top}X^{\top}\mathbf{1}_{m}+\epsilon\mathbf{1}_{m})\cdot XM\cdot\text{diag}(\omega^{(l)}),$
@@
B.4 Proof for Theorem 4.4 and Related Discussions
We first provide the proof for Theorem 4.4 based on Section A.1.
@@
On the other hand, semi-orthogonal $U_{B,\text{Q}}^{(l)},U_{B,\text{K}}^{(l)}$ are also possible. This is because of the common practice in pruning may select different different rows of Q,K matrices, which suggests multiplying Q,K matrices with row-wise partial permutations.
@@
B.5 How Potential Attacks in Section 4.5 Recovers the Output of Transformers
We provide an illustration based on Section A.1, Section A.1 and Section A.1. Previous parts have demonstrated how the manipulations on inputs pass through self-attention. Hence, it suffices to show how the manipulations pass through FFN and yield an output identical to model A after the language model head.
@@
By Section A.1 and Section A.1, the manipulation is propagated through transformer layers. At the language model head, we denote $W_{B,\text{lm}}=c^{-1}W_{A,\text{lm}}PD$ and abuse $X^{\prime}=cXPD$ as the manipulated input. Then,
@@
Appendix C Mapping of Model abbreviations
LABEL:tab:model_abbreviations details the mapping between the abbreviations used throughout the paper, their respective base models, and their full model names as cataloged on the Hugging Face Hub.
@@ Appendix C Mapping of Model abbreviations
Table 4: Mapping of model abbreviations to their full Hugging Face model names.
Abbreviation Base Model Full Model Name
Vicuna Llama2-7B vicuna-7b-v1.5
Finance Llama2-7B llama-2-7b-finance
Selfrag Llama2-7B selfrag_llama2_7b
32K Llama2-7B LLaMA-2-7B-32K
Wizard Llama2-7B WizardMath-7B-V1.0
Guanaco Llama2-13B llama-2-7b-guanaco
Vicuna Llama2-13B vicuna-13b-v1.5
Hermes Llama2-13B Nous-Hermes-Llama2-13b
Estopia Llama2-13B LLaMA2-13B-Estopia
Firefly Llama2-13B firefly-llama2-13b
Llemma Llama2-7B llemma_7b
Code Llama2-7B CodeLlama-7b-hf
Python Llama2-7B CodeLlama-7b-Python-hf
Code Gemma-2B codegemma-2b
Code Gemma-7B codegemma-7b
Math Qwen2.5-7B Qwen2.5-Math-7B
Coder Qwen2.5-7B Qwen2.5-Coder-7B
Math Qwen2-7B Qwen2-Math-7B
Code Llama2-70B CodeLlama-70b-hf
Python Llama2-70B CodeLlama-70b-Python-hf
Mixtral Mistral-7B Nous-Hermes-2-Mixtral-8x7B-DPO
MoE v2 Llama3-8B LLaMA-MoE-v2-3_8B-2_8-sft
MoE4 Llama2-7B LLaMA-MoE-v1-3_5B-4_16
MoE 3B Llama2-7B LLaMA-MoE-v1-3_0B-2_16
MoE2 Llama2-7B LLaMA-MoE-v1-3_5B-2_8
MoE3B-SFT Llama2-7B LLaMA-MoE-v1-3_0B-2_16-sft
MoE2-SFT Llama2-7B LLaMA-MoE-v1-3_5B-2_8-sft
MoE4-SFT Llama2-7B LLaMA-MoE-v1-3_5B-4_16-sft
Qwen1.5 MoE Qwen-1.8B Qwen1.5-MoE-A2.7B
Minicpm MoE Minicpm-2B MiniCPM-MoE-8x2B
LLaVA Llama2-7B llava-v1.5-7b
Video Llama2-7B Video-LLaVA-7B-hf
VL Qwen2-7B Qwen2-VL-7B-Instruct
Audio Qwen-7B Qwen-Audio
Audio2 Qwen-7B Qwen2-Audio-7B
VL Qwen-7B Qwen-VL
VL Qwen2.5-7B Qwen2.5-VL-7B-Instruct
VL Qwen2.5-3B Qwen2.5-VL-3B-Instruct
Next Llama3-8B llama3-llava-next-8b-hf
LLaVA Llama2-13B llava-v1.5-13b
RLHF Open-llama3B hh_rlhf_rm_open_llama_3b
Reason Qwen2.5-7B Nemotron-Research-Reasoning-Qwen-1.5B
Zero Qwen2.5-1.5B Open-Reasoner-Zero-1.5B
DPO Mixtral Nous-Hermes-2-Mixtral-8x7B-DPO
DPO Mistral-7B Nous-Hermes-2-Mistral-7B-DPO
Dolphin Mistral-7B dolphin-2.6-mistral-7b-dpo
DPO Minicpm-2B MiniCPM-2B-dpo-bf16
GRPO Qwen3-4B Qwen3_Medical_GRPO
RLHF Chatglm-6B chatglm-fitness-RLHF
DPO Llama3-8B LLaMA3-iterative-DPO-final
Minitron-Depth Llama-3-8B Llama-3.1-Minitron-4B-Depth-Base
Minitron-Width Llama-3-8B Llama-3.1-Minitron-4B-Width-Base
Llama3-1B Llama3-8B Llama-3.2-1B
Llama3-3B Llama3-8B Llama-3.2-3B
Sheared 2.7B-P Llama2-7B Sheared-LLaMA-2.7B-Pruned
Sheared 2.7B-S Llama2-7B Sheared-LLaMA-2.7B-ShareGPT
Sheared 2.7B Llama2-7B Sheared-LLaMA-2.7B
Sheared 1.3B-P Llama2-7B Sheared-LLaMA-1.3B-Pruned
Sheared 1.3B Llama2-7B Sheared-LLaMA-1.3B
Sheared 1.3B-S Llama2-7B Sheared-LLaMA-1.3B-ShareGPT

### v2

@@ 6 Experiments
Baselines. We compare our method against two advanced baselines: the weight-based method HuRef (zeng2024huref), and the representation-based method REEF (zhang2025reef).
@@ 6.3 Robustness Verification on 60 Offspring Models
Table 2: Absolute Z-scores ($\uparrow$) of positive samples for HuRef, REEF, and our method. The full names of abbreviations and related model details are provided in Appendix G.
@@ 6.4 Overall Evaluation
Intrinsic Fingerprint ||Z|| $\uparrow$ 1.535 1.193 1.408 1.532 1.542 1.141 1.392
AUC $\uparrow$ 1.000 0.896 0.969 1.000 1.000 0.876 0.957
pAUC $\uparrow$ 1.000 0.422 0.800 1.000 1.000 0.400 0.770
TPR@1%FPR $\uparrow$ 1.000 0.300 0.800 1.000 1.000 0.400 0.750
PCS ||Z|| $\uparrow$ 73.786 74.650 0.727 14.950 163.399 17.226 57.457
AUC $\uparrow$ 0.958 0.959 0.791 0.802 0.984 0.666 0.860
pAUC $\uparrow$ 0.656 0.600 0.078 0.500 0.900 0.100 0.472
TPR@1%FPR $\uparrow$ 0.500 0.600 0.000 0.500 0.900 0.100 0.433
@@
To synthesize our findings, we conducted a comprehensive evaluation against HuRef (ICS) and REEF using the full dataset of 60 positive and 90 negative model pairs. Additionally, we incorporated PCS (from HuRef) and Intrinsic Fingerprint (yoon2025intrinsic) as supplementary baselines to broaden the comparative analysis.
@@ 7 Conclusions
Appendix A Ethics Statement
The primary motivation for this work is to protect the intellectual property of large language models (LLMs), thereby promoting fairness and accountability within the AI community. We acknowledge the potential societal impact of such a technology, particularly the risk of false accusations of model theft. Therefore, a core design principle of our method is to achieve an exceptionally low false-positive rate, as empirically demonstrated in our experiments. By establishing a high-fidelity verification system, we aim to foster a more transparent and trustworthy open-source ecosystem. All models used in this study are publicly available, and our research does not involve any private or sensitive user data.
Appendix B Reproducibility Statement
To ensure the reproducibility of our results, we have included the complete source code in the supplementary material. We also provided a detailed description of our methodology in Section 5, including a step-by-step algorithm. All models used for evaluation are publicly available from sources such as the Hugging Face Hub, and a comprehensive list mapping our model abbreviations to their full names is included in Appendix G. A public code repository will be made available upon publication.
Appendix C The Use of Large Language Models
The core research and analysis presented in this manuscript were conducted without the use of Large Language Models (LLMs). An LLM was utilized exclusively to improve the language and clarity of the text.
Appendix D Definitions and Notations
D.1 LLM Architecture
@@
$\mathrm{RMSNorm}(X,\omega_{A}^{(l)},\epsilon_{A})=\operatorname{diag}^{-\frac{1}{2}}\!\Big((X\odot X)\mathbf{1}_{n}+\epsilon_{A}\mathbf{1}_{m}\Big)\,X\,\operatorname{diag}(\omega_{A}^{(l)}).$ (6)
Definition D.2 (Self-Attention with RoPE).
@@
Here $x_{i}^{\top}\in\mathbb{R}^{n}$ is the i-th row of $X$, $\theta$ is the frequency base for RoPE and $R_{\theta}$ is the rotation matrix (su2024roformer).
RoPE’s definition in Section D.1 is slightly different from their implementations, but still an equivalent version to su2024roformer.
Definition D.3 (Feed-forward Networks).
@@
We abuse the notations in Section D.1 by denoting $\text{RMSNorm}_{\text{attn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})$ as the attention norm and $\text{RMSNorm}_{\text{ffn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{ffn}}^{(l)},\epsilon_{A})$ the ffn norm at layer $l$ of model A. Then, the definition of a Transformer block at layer $l$ of model A can be summartized as follows.
Definition D.4 (Transformer Block).
@@
D.2 Central Kernel Alignment (CKA)
@@
where $\text{HSIC}(K_{X_{1}},K_{X_{2}})=\frac{1}{(m-1)^{2}}\text{tr}(K_{X_{1}}H_{m}K_{X_{2}}H_{m})$, $(K_{X_{1}})_{ij}=k((X_{1})_{i},(X_{1})_{j})$, $(K_{X_{2}})_{ij}=k((X_{2})_{i},(X_{2})_{j})$ are kernel matrices with the kernel function $k(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}_{+}$, and $H_{m}=I_{m}-\frac{1}{m}\mathrm{1_{m}}\mathrm{1_{m}}^{\top}$ is the centering matrix. The linear kernels give $K_{X}=XX^{\top}$ for a matrix $X$. To reduce the finite-sample bias of Eq. (6), we use the unbiased HSIC estimator (song2007supervised). Let $\tilde{K}_{X_{i}}$ be $K_{X_{i}}$ with its diagonal set to zero, i.e., $(\tilde{K}_{X_{i}})_{ii}=0$. With $\mathrm{1}_{m}$ the all-ones vector of length $m$, the unbiased HSIC is
@@
Appendix E Proofs and Discussions
E.1 Proof for Theorem 3.1
We give the proof based on the definition of CKA with linear kernels in Appendix D.2. Given input matrices $X_{1},X_{2}$, orthogonal matrices $U_{1},U_{2}$ and constants $c_{1},c_{2}$, the kernel functions yield
@@ Appendix E Proofs and Discussions
E.3 Proof for Theorem 4.3 and Discussions on Other Input Manipulations
We first show a proof for the claims in Theorem 4.3. The core lies in preserving the diagonal structure of RMSNorm parameters in Section D.1. Given the input $X$ to layer $l$, the manipulations change it to $cXPD$ and fed the manipulated input into the attention norm of model B. In order to recover the manipulations on inputs after RMSNorm, i.e. for any manipulation $T$ in the input $X$ (an example is $T(X)=cXPD$),
@@
$=\text{diag}^{-\frac{1}{2}}\Big(\big((cXPD)\odot(cXPD)\big)\mathbf{1}_{n}+\epsilon_{B}\mathbf{1}_{m}\Big)\cdot cXPD\cdot\text{diag}(\omega_{B,\text{attn}}^{(l)})$
$=\text{diag}^{-\frac{1}{2}}\Big(c^{2}(X\odot X)\mathbf{1}_{n}+c^{2}\epsilon_{A}\mathbf{1}_{m}\Big)\cdot cXPD\cdot cD^{\top}P^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})PD$
$=c\cdot\text{diag}^{-\frac{1}{2}}\Big((X\odot X)\mathbf{1}_{n}+\epsilon_{A}\mathbf{1}_{m}\Big)\cdot X\cdot\text{diag}(\omega_{A,\text{attn}}^{(l)})\cdot PD$
@@
However, since $(X\odot X)\mathbf{1}_{n}=(XU\odot XU)\mathbf{1}_{n}$, we have
$\text{RMSNorm}(XU,\omega,\epsilon)=\text{diag}^{-\frac{1}{2}}\Big((X\odot X)\mathbf{1}_{n}+\epsilon\mathbf{1}_{m}\Big)\cdot XU\cdot\text{diag}(\omega^{(l)}).$
@@
$\text{RMSNorm}(XM,\omega^{(l)},\epsilon)=\text{diag}^{-\frac{1}{2}}\Big((XM\odot XM)\mathbf{1}_{n}+\epsilon\mathbf{1}_{m}\big)\cdot XM\cdot\text{diag}(\omega^{(l)}),$
@@
E.4 Proof for Theorem 4.4 and Related Discussions
We first provide the proof for Theorem 4.4 based on Section D.1.
@@
On the other hand, semi-orthogonal $U_{B,\text{Q}}^{(l)},U_{B,\text{K}}^{(l)}$ are also possible. This is because of the common practice in pruning may select different rows of Q,K matrices, which suggests multiplying Q,K matrices with row-wise partial permutations.
@@
E.5 How Potential Attacks in Section 4.5 Recovers the Output of Transformers
We provide an illustration based on Section D.1, Section D.1 and Section D.1. Previous parts have demonstrated how the manipulations on inputs pass through self-attention. Hence, it suffices to show how the manipulations pass through FFN and yield an output identical to model A after the language model head.
@@
By Section D.1 and Section D.1, the manipulation is propagated through transformer layers. At the language model head, we denote $W_{B,\text{lm}}=c^{-1}W_{A,\text{lm}}PD$ and abuse $X^{\prime}=cXPD$ as the manipulated input. Then,
@@
Appendix F Ablation Studies
F.1 Number of Overlapping Vocabulary Tokens
Although Algorithm 1 uses overlapping tokens to recover the signature matrices and permutations, the detection performance of AWM does not heavily depend on the amount of vocabulary overlap. In fact, AWM remains effective even when only a small number of tokens ( 100 tokens) are shared between the two vocabularies. To quantify this, we conduct an ablation study on the number of overlapping tokens used in Algorithm 1. Specifically, for each scenario in Table 2 (SFT, Continual Pretraining, Upcycling, Multi-Modal, RL, and Pruning), we compute the average absolute Z-score under different numbers of overlapping vocabulary tokens and report the results in Figure 3.
Figure 3: Ablation studies on the number of overlapping vocabulary tokens vs AWM’s average absolute Z-score in Table 2. AWM remain effective even when there are only 100 overlapping tokens used in Algorithm 1.
F.2 CKA Ablations
We further conduct an ablation study on the design of CKA in Algorithm 1. In AWM, we use the unbiased version with linear kernel for computation efficiency and accuracy. Here we investigate into more variants of CKA, and summarize the results in Table 4. It can be seen that the choice of unbiasedness is crucial to the robustness of our method. Meanwhile, although RBF kernels show stronger performance in Table 4, the choice of linear kernel has already yield strong performance, which we choose for better computational efficiency.
CKA Kernel Linear RBF
CKA Biasedness Unbiased Biased Unbiased Biased
SFT 356.0223 18.3099 527.3949 11.5843
Continual Pretrain 217.5003 11.2954 320.9173 7.0701
Upcycling 291.6191 15.1994 432.1476 9.6776
Multi Modal 336.6757 17.3766 498.8918 11.0056
RL 357.5001 18.3850 529.6448 11.6330
Pruning 268.9175 13.8985 394.8391 8.7331
Table 4: AWM’s average absolute Z-scores under CKA with different kernels and biasedness.
Appendix G Model Details
LABEL:tab:model_abbreviations_with_corpus details the specifications of the offspring models analyzed in our study, mapping abbreviations to their base models and training datasets. To facilitate reproducibility, each entry in the "Full Model Name" column serves as a direct link to the official model checkpoint hosted on the Hugging Face Hub.
@@ Appendix G Model Details
Table 5: Mapping of model abbreviations to their full model names, corresponding Hugging Face checkpoints url, base models, and relevant training corpus information.
Abbreviation Base Model Full Model Name Train Corpus
Vicuna Llama2-7B vicuna-7b-v1.5 ShareGPT
Selfrag Llama2-7B selfrag_llama2_7b Self-RAG Data
32K Llama2-7B LLaMA-2-7B-32K Book, ArXiv, etc.
Wizard Llama2-7B WizardMath-7B-V1.0 GSM8k
Guanaco Llama2-13B llama-2-7b-guanaco OpenAssistant
Vicuna Llama2-13B vicuna-13b-v1.5 ShareGPT
Hermes Llama2-13B Nous-Hermes-Llama2-13b GPTeacher, WizardLM, etc.
Estopia Llama2-13B LLaMA2-13B-Estopia EstopiaV9/V13, Tiefighter, etc.
Finance Llama2-7B llama-2-7b-finance Financial Dataset
Firefly Llama2-13B firefly-llama2-13b CLUE, ThucNews, etc.
Llemma Llama2-7B llemma_7b ArXiv, OpenWebMath, etc.
Code Llama2-7B CodeLlama-7b-hf Deduped code, Natural language
Python Llama2-7B CodeLlama-7b-Python-hf Python code
Code Gemma-2B codegemma-2b Math, Synthetic code, etc.
Code Gemma-7B codegemma-7b Code, Natural language
Math Qwen2.5-7B Qwen2.5-Math-7B Web, Books, etc.
Coder Qwen2.5-7B Qwen2.5-Coder-7B Source Code, Synthetic data, etc.
Math Qwen2-7B Qwen2-Math-7B Math Data
Code Llama2-70B CodeLlama-70b-hf Deduped code, Natural language
Python Llama2-70B CodeLlama-70b-Python-hf Python code
Mixtral Mistral-7B Nous-Hermes-2-Mixtral-8x7B-DPO GPT-4 Data, Open datasets
MoE v2 Llama3-8B LLaMA-MoE-v2-3_8B-2_8-sft SFT Data
MoE4 Llama2-7B LLaMA-MoE-v1-3_5B-4_16 SlimPajama
MoE 3B Llama2-7B LLaMA-MoE-v1-3_0B-2_16 SlimPajama
MoE2 Llama2-7B LLaMA-MoE-v1-3_5B-2_8 SlimPajama
MoE3B-SFT Llama2-7B LLaMA-MoE-v1-3_0B-2_16-sft SlimPajama, SFT Data
MoE2-SFT Llama2-7B LLaMA-MoE-v1-3_5B-2_8-sft SlimPajama, SFT Data
MoE4-SFT Llama2-7B LLaMA-MoE-v1-3_5B-4_16-sft SlimPajama, SFT Data
Qwen1.5 MoE Qwen-1.8B Qwen1.5-MoE-A2.7B Qwen Base Corpus
Minicpm MoE Minicpm-2B MiniCPM-MoE-8x2B MiniCPM Data
LLaVA Llama2-7B llava-v1.5-7b LAION, GPT instructions, etc.
Video Llama2-7B Video-LLaVA-7B-hf Caption, QA
VL Qwen2-7B Qwen2-VL-7B-Instruct Image-text, OCR, etc.
Audio Qwen-7B Qwen-Audio Speech, Sound, etc.
Audio2 Qwen-7B Qwen2-Audio-7B Audio-text, Voice Chat
VL Qwen-7B Qwen-VL Image-text, OCR, etc.
VL Qwen2.5-7B Qwen2.5-VL-7B-Instruct Visual recognition, Document parsing, etc.
VL Qwen2.5-3B Qwen2.5-VL-3B-Instruct Visual recognition, Document parsing, etc.
Next Llama3-8B llama3-llava-next-8b-hf LLaVA-NeXT Data
LLaVA Llama2-13B llava-v1.5-13b LAION, GPT instructions, etc.
RLHF Open-llama3B hh_rlhf_rm_open_llama_3b Anthropic HH-RLHF
Reason Qwen2.5-7B Nemotron-Research-Reasoning-Qwen-1.5B Math, Code, etc.
Zero Qwen2.5-1.5B Open-Reasoner-Zero-1.5B AIME 2024, MATH500, etc.
DPO Mixtral Nous-Hermes-2-Mixtral-8x7B-DPO GPT-4 Data, Preference pairs
DPO Mistral-7B Nous-Hermes-2-Mistral-7B-DPO GPT-4 Data, Preference pairs
Dolphin Mistral-7B dolphin-2.6-mistral-7b-dpo UltraFeedback, Magicoder, etc.
DPO Minicpm-2B MiniCPM-2B-dpo-bf16 ShareGPT, UltraChat, etc.
GRPO Qwen3-4B Qwen3_Medical_GRPO Medical dataset
RLHF Chatglm-6B chatglm-fitness-RLHF SFT, Reward Model, etc.
DPO Llama3-8B LLaMA3-iterative-DPO-final UltraFeedback, Preference sets
Minitron-Depth Llama-3-8B Llama-3.1-Minitron-4B-Depth-Base Nemotron-4 15B corpus
Minitron-Width Llama-3-8B Llama-3.1-Minitron-4B-Width-Base Nemotron-4 15B corpus
Sheared 2.7B-P Llama2-7B Sheared-LLaMA-2.7B-Pruned RedPajama
Sheared 2.7B-S Llama2-7B Sheared-LLaMA-2.7B-ShareGPT ShareGPT
Sheared 2.7B Llama2-7B Sheared-LLaMA-2.7B RedPajama
Sheared 1.3B-P Llama2-7B Sheared-LLaMA-1.3B-Pruned RedPajama
Sheared 1.3B Llama2-7B Sheared-LLaMA-1.3B RedPajama
Sheared 1.3B-S Llama2-7B Sheared-LLaMA-1.3B-ShareGPT ShareGPT
Llama3-1B Llama3-8B Llama-3.2-1B Llama3-8B logits, Safety data
Llama3-3B Llama3-8B Llama-3.2-3B Llama3-8B logits, Safety data
Appendix H Implementation Details
We employ the Linear Kernel ($k(X,Y)=XY^{\top}$) for Centered Kernel Alignment (CKA) due to its computational efficiency. To mitigate the finite-sample bias inherent in standard HSIC estimations, we utilize the Unbiased CKA (UCKA) estimator. As for module selection, our method operates on two specific sets of weights: first, we utilize the intersection of the word embeddings ($W_{emb}$) to solve the Linear Assignment Problem (LAP), allowing us to accurately recover the permutation ($P$) and signature ($D$) matrices; second, we compute the final fingerprinting scores using the Query ($W_{Q}$) and Key ($W_{K}$) weights, as their transformations are strictly constrained.
To address structural discrepancies such as differing layer counts, we identify the optimal layer correspondence by maximizing the total similarity; specifically, we solve the assignment problem on a cost matrix constructed from the pairwise UCKA scores of $W_{Q}$ and $W_{K}$ between all source and target layers.
Appendix I Empirical Validation of Robustness against Weight Manipulations
In Section 4, we theoretically analyze potential weight manipulations, including constant scaling, signature matrix multiplication, permutations, and orthogonal transformations. We now provide empirical evidence to support the analysis. Specifically, we apply these manipulations to five representative models, Llama-2-7B, Qwen2-7B, Mistral-7B, Llama-3-8B, and Gemma-7B, and evaluate AWM’s robustness under each setting.
As shown in Table 6, AWM achieves a 100% detection rate across all tested model–manipulation pairs. These results not only align with our theoretical derivations, but also validate the design of AWM, including LAP and UCKA.
Table 6: AWM-detected similarity scores under the weight manipulations in Section 4.
Manipulation / Model Llama-2-7B Qwen2-7B Mistral-7B Llama-3-8B Gemma-7B
Permutation 100% 100% 100% 100% 100%
Signature 100% 100% 100% 100% 100%
Constant Scaling 100% 100% 100% 100% 100%
Orthogonal Trans. 100% 100% 100% 100% 100%